a number when divided by 2 leaves a remainder 3 . when it is divided by 3, the remainder is 2. what is the number
Answer by Theo(13342)

a number divided by 2 cannot leave a remainder of 3.

example:

1/2 = 0 with a remainder of 1.
2/2 = 1 with a remainder of 0.
3/2 = 1 with a remainder of 1.
4/2 = 2 with a remainder of 0.

0/2 = 0 with a remainder of 0.
-1/2 = 0 with a remainder of -1.
-2/2 = -1 with a remainder of 0.
-3/2 = -1 with a remainder of -1.
-4/2 = -2 with a remainder of 0.

that's my take on this problem.

if i'm correct, there is no number divided by 2 that leaves a remainder of 3.